A mainland population of 1000 blue iguanas have the neutral alleles A1 and A2 with frequencies of A1 = 0.50 and A2 = 0.50. A new population of blue iguanas is founded on Lizard Island with a population size of 10 males and 15 females. The neutral alleles A1 and A2 are present in the population with frequencies of A1 = 0.30 and A2 = 0.70.

Migrants begin moving from the mainland to Lizard Island at a rate of 1 migrant per generation.

1. Calculate the change in frequency (?p) of the A1 allele after one generation of migration.

2. What will the new frequency (p?) of A1 be in the island population after one generation of migration ?
Answer #1

Mainland population:

No. of blue iguanas = 1000

Allele frequency of A1 = 0.50

Lizard island population:

No. of blue iguanas = 25

Allele frequency of A1 = 0.30

Now, given 1 migrant per generation from mainlaind to island

rate of migration = 1/26 = 0.038 (26 because after one generation net population is 25+1=26)

Ans 2. New frequency of island population (p') = (Allele frequency of residents)(1- migration rate) + (migration rate)(allele frequency of migrating population)

p' = (0.3)(1-0.038) + (0.038)(0.5)

p' = 0.2886 + 0.019 = 0.3076

Ans 1. Change in frequency (delta p) = New allele frequency on island - old allele frequency on island

= 0.3076 - 0.30 = 0.0076